please help solve this equation
3x-7y+5(x+3y)-4(y-2x)

Found 3 solutions by ewatrrr, MathLover1, ezan:
Answer by ewatrrr(24785)

Removing Parenthesis and combining Like Terms...
3x-7y+5(x+3y)-4(y-2x)= 3x-7y+5x + 15y - 4y + 8x = 16x+4y

Answer by MathLover1(20850)
Answer by ezan(1)

3x-7y+5(x+3y)-4(y-2x)= 3x-7y+5x + 15y - 4y + 8x = 16x+4y